Each shot of the laser gun most favored by Rosa the Closer, the intrepid vigilante of the lawless 22nd century, is powered by the discharge of a 1.37 F1.37 F capacitor charged to 77.1 kV77.1 kV. Rosa rightly reckons that she can enhance the effect of each laser pulse by increasing the electric potential energy of the charged capacitor. She could do this by replacing the capacitor's filling, whose dielectric constant is 463463, with one possessing a dielectric constant of 961961.

Find the electric potential energy of the original capacitor when it is charged.

original capacitor potential energy:

JJ

Calculate the electric potential energy of the upgraded capacitor when it is charged.

upgraded capacitor potential energy:
Answer #1

Part A.

Potential energy stored in capacitor is given by:

PE = (1/2)*C*V^2

C = Capacitance of capacitor when 1st dielectric is inserted = 1.37 F

V = Potential difference across plates = 77.1 kV = 77100 V

So,

PE = (1/2)*1.37*77100^2

PE = 4.07*10^9 J

Part B.

Now when initial dielectric (k1) is removed and another dielectric (k2) is inserted, then

C1 = k2*C/k1 = (k2/k1)*C

Since battery is still connected, So voltage will remain same. V1 = V

So now electric potential energy of capacitor will be:

PE1 = (1/2)*C1*V1^2

PE1 = (1/2)*(k2/k1)*C*V^2

PE1 = (1/2)*(961/463)*1.37*(77100)^2

PE1 = 8.45*10^9 J

Let me know if you've any query.